OpenStudy (anonymous):

Find the antideravative for the given function: f(x) = sinx/(cosx)^2

OpenStudy (turingtest):

\[u=\cos x\]what then is \(du\) ?

OpenStudy (anonymous):

(cosx)^2?

OpenStudy (anonymous):

Oh wait, du would be, -sinx

OpenStudy (turingtest):

\[du=-\sin xdx\]the dx is actually very important, but yes

OpenStudy (anonymous):

Okay

OpenStudy (turingtest):

the integral you have is\[\int{\sin xdx\over\cos^2x}\]now substitute\[u=\cos x\]\[du=-\sin xdx\implies\sin xdx=-du\]into the integral what do you get?

OpenStudy (anonymous):

I'm sorry, I'm actually confused.

OpenStudy (anonymous):

It's just that we have been introduced to this method yet, so I'm not sure what you're doing.

OpenStudy (anonymous):

have not been *

OpenStudy (turingtest):

we are making use of the fact that the numerator is almost the derivative of the function in the denominator do yo agree that\[du=-\sin xdx\]?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

so then can you see that\[\sin xdx=-du\] ?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

so we can substitute in\[\int{\sin xdx\over\cos^2x}\]\[\cos x=u\]\[\sin xdx=-du\]and we get\[\int\frac{-du}{u^2}=-\int u^{-2}du\]which you should be able to integrate now

OpenStudy (anonymous):

so the answer is -1/cosx + c?

OpenStudy (turingtest):

i thinnk the -

OpenStudy (turingtest):

...goes away

OpenStudy (anonymous):

Okay, got it.Thank you! :)

OpenStudy (turingtest):

welcome :)